Industry demand Qd is given by Qd= 26−2P (where P is industry price), and the marginal cost MC of industry output Qs is given by MC = Qs−2?
(a) What is the industry price and quantity if there are a large number of buyers
and sellers in the industry.
(b) What is meant by consumer surplus. What is the level of consumer surplus
in this industry.
1
Expert's answer
2017-03-22T11:01:06-0400
Answer:
A) P=MR=MC
Qs=P+2
26-2P=P+2
P=8
Qd=Qs=10

b) Consumer surplus  is calculated by analyzing the difference between what consumers are willing and able to pay for a good or service relative to its market price, or what they actually do spend on the good or service. A consumer surplus occurs when the consumer is willing to pay more for a given product than the current market price.

CS=0.5*(10-0)*(13- 8)=25